hey guy, i need your help in this question , thank you soooo much!

More Info Variable Definitions Variable ??? College Female Age Ntheast Midwest South West Definition Average hourly earnings (in 1998 dollars) Binary variable (1 if college, 0 if high school) Binary variable (1 if female, 0 if male). Age (in years) Binary variable (1 if RegionNortheast, 0 otherwise) Binary variable (1 if Region Midwest, 0 otherwise) Binary variable (1 if Region South, 0 otherwise) Binary variable (1 if RegionWest, 0 otherwise) Print Done
Answer #1

Adjusted R2 = 1 - [(1 - R2) x (n - 1)} / (n - p - 1)] where

n: Number of observations = 4,701 and

p = Number of independent variables = 6

Adjusted R2 = 1 - [(1 - R2) x (4,701 - 1)} / (4,701 - 6 - 1)] = 1 - [(1 - R2) x 4,000} / 4,694] = 1 - [0.8522 x (1 - R2)]

Model 1: R2 = 0.143

Adjusted R2 = 1 - [0.8522 x (1 - 0.143)] = 1 - (0.8522 x 0.857) = 1 - 0.730 = 0.270

Model 2: R2 = 0.168

Adjusted R2 = 1 - [0.8522 x (1 - 0.168)] = 1 - (0.8522 x 0.832) = 1 - 0.709 = 0.291

Model 3: R2 = 0.143

Adjusted R2 = 1 - [0.8522 x (1 - 0.124)] = 1 - (0.8522 x 0.876) = 1 - 0.747 = 0.253